# COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Gravitational Lenses in DECaLS

### 17 Jan 2024 v2

@@ Abstract
We compute parametric measurements of the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically-selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $\sim 0.2$ to $\sim 1.0$ using only two parameters: the Einstein radius and the brightest-cluster-galaxy (BCG) redshift. We constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions (SEDs) with aperture photometry from the Dark Energy Camera Legacy Survey (DECaLS) in the $grz$ Dark Energy Camera (DECam) filters. We find that the BCG redshift, enclosed total mass, and enclosed luminosity are strongly correlated and well described by a planar relationship in 3D space. We also find that the enclosed total mass and stellar mass are correlated with a logarithmic slope of $0.443\pm 0.035$, and the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope of $-0.563\pm 0.035$. The correlations described here can be used to validate strong lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time (LSST)—in which an algorithmic treatment of lensing systems will be needed due to the sheer volume of data these surveys will produce.
@@ 1 Introduction
Strong gravitational lensing is a rare phenomenon in the universe in which an intervening object with sufficient surface mass density bends light from a background source relative to the observer such that multiple images of the source are formed. For cases in which the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arclike shape (e.g., Lynds & Petrosian 1986; Soucail et al. 1987; Kneib & Natarajan 2011; Meneghetti et al. 2013; Rivera-Thorsen et al. 2017; Shajib et al. 2022; Zhang et al. 2023). The discovery of new gravitational lenses of this type is accelerating rapidly, particularly with the help of machine learning applied to finding galaxy-galaxy strong lenses in large optical-imaging surveys (e.g., Huang et al. 2021; Rojas et al. 2022; Zaborowski et al. 2022). Coupled with access to a flood of new data from large surveys such as Rubin/LSST in the coming years, the population of candidate lenses both in the bulk and at the margins will only grow. Contemporary methodologies often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (Diehl et al., 2017); such inspection has been successful in locating new systems thus far (Rojas et al., 2023). However, with thousands to millions of lines of sight to choose from in an ever-growing plethora of imaging data, statistical measures and correlations in existing lensing systems may help hone candidate-lens samples by quantifying instances of strong lensing in tandem with traditional morphological identification. This would produce a purer sample and help streamline the ability to identify targets. In this paper, we analyze a sample of 177 strong gravitational lenses identified via visual inspection by the COOL-LAMPS collaboration in DECaLS Legacy Survey Data Release 8 (LS DR8, Dey et al. 2019) images, and we investigate the scaling relations between the BCG redshift, Einstein-radius-enclosed (enclosed) total mass, enclosed luminosity, enclosed stellar mass, and enclosed stellar-to-total mass fraction therein. The sample we use here consists of a refined subsample of all strong-lensing candidates identified by the COOL-LAMPS team, and we consider them unambiguous and robust as informed by their visual morphology—despite a lack of definite spectroscopic confirmation.
@@ 2 Methodology
Galaxies and galaxy clusters host their mass in a variety of forms. Theory and observations both place strong constraints on observed stellar luminosity as a function of mass (e.g., Kuiper 1938; Presotto et al. 2014; Wang & Zhong 2018), hot intracluster gas contributes significantly to the total mass in galaxy clusters (e.g., Kellogg et al. 1972; Forman et al. 1972; Markevitch & Vikhlinin 2007), and both galaxies and clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al. 1996; Bryan & Norman 1998; Hradecky et al. 2000). However, robustly measuring these mass components on a per-system basis is time and resource intensive—which is unsustainable for large samples. Conveniently, the existence of lensing prescribes a specific amount of mass that must exist to create the observed lensing in the first place.
@@
Remolina González et al. (2020) have shown that knowing only the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$), the core mass interior to the Einstein radius of a galaxy cluster may be measured with minimal scatter on order of $\approx 10\%$. With these three variables, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius. Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties as well. By directly integrating over pixels inside this aperture in flux-calibrated images, the photometry can be automatically constrained and utilized by an SED modeller to obtain measurements of enclosed luminosity and stellar mass.
@@
The COOL-LAMPS collaboration has constructed a large dataset of strong lensing candidates using a simple human-led ranking system (e.g., Khullar et al. 2021). From this, we have collated a sample of 177 galaxy clusters and groups that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed arc is roughly circular with its projected center located at or very near to the BCG. Any systems with complex source-galaxy morphology are less likely to be well-described by the simple symmetric dark-matter profiles that inform the parametric mass estimates described in Remolina González et al. (2020). Such complex systems were intentionally excluded from the sample considered in this work.
@@
Figure 1: A subset of nine strong gravitational lenses used in this work. The lenses are arranged by increasing redshift from the top to the bottom of the figure, and they are arranged by increasing Einstein radius from the left to the right of the figure. Red circles denote positional constraints used to derive the Einstein radius in each of the systems. White circles represent the Einstein aperture for each system. A black ‘x’ at the center of each subplot marks the exact coordinates of the BCG. The lens redshifts and Einstein radii for each system are also shown in each panel. RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using the $z$, $r$, and $g$ bands with custom color scaling respectively.
@@
Figure 2: Observed-frame spectral energy distributions (in maggies) for the nine systems shown in Figure 1 after SED fitting with Prospector. The top plot of each subplot contains the best-fit spectrum (red line), 1024 spectra drawn from their respective posterior distribution (grey lines), observed photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam, Flaugher et al. 2015) $g$, $r$, and $z$ filters from left to right respectively (in arbitrary units). The bottom plot of each subplot shows the residual between the best-fit spectra and the observed photometry normalized by the standard deviation of the observed photometry.
@@ 2.1 Einstein Radius
We defined the Einstein radius for each lensing system as the radius of a circle centered on the BCG (Remolina González et al., 2020) which minimizes the total angular separation between each of the three most-readily-identifiable bright “clumps” in the tangentially-lensed arcs and the perimeter of said circle. Henceforth, we refer to this circular aperture centered on the BCG with radius equal to the Einstein radius as the “Einstein aperture”. Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially-lensed arc respectively. If three distinct peaks were unable to be found along the tangential arcs, the main arc was simply traced. We bootstrapped the three tangential-arc positional constraints 1000 times, and we adopted the mean and standard deviation for the resulting distribution of Einstein radii as the Einstein radius and its error respectively for each system. A descriptive visualization sampling the range of Einstein radii and lens redshifts considered in this work is shown in Figure 1.
@@ 2.2 Lens Redshift
BCGs are sui generis red-sequence cluster galaxies which can also be used as proxies to infer properties of the cluster-scale dark matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al. 2014). Since we assume that the lensing systems in this work can be well-described by a symmetric dark-matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale halo. The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we query the Sloan Digital Sky Survey Data Release 15 (SDSS DR15, Aguado et al. 2019; Kollmeier et al. 2019) in tandem with the Legacy Survey Data Release 9 (LS DR9, Dey et al. 2019; Duncan 2022) for each system in this work to obtain BCG redshifts. If the centroid coordinates for each BCG corresponded to a warning-free spectroscopic redshift in SDSS DR15, we adopted that redshift as the lens redshift for that system along with its associated error. If a given BCG lacked any corresponding value in SDSS DR15, as is the case for the higher redshift clusters in this work and/or clusters located in SDSS-non-imaged areas, we adopted the LS DR9 photometric redshift (Zhou et al., 2023) as the lens redshift and its associated error instead. Summarized in Table 2, we obtained spectroscopic redshifts for 41 systems, and photometric redshifts for the remaining 136.
Table 1: Free Parameters used in SED fitting with Prospector
Parameter Description Priors
$\mathrm{log}(M_{tot}/M_{\odot})$ Total stellar mass formed in dex solar masses. Top Hat: [8.0, 14.0]
$\mathrm{log}(Z/Z_{\odot})$ Stellar metallicity in dex solar metallicity. Top Hat: [-1.0, 0.2]
$\lambda_{2}$ Diffuse dust optical depth. Top Hat: [0.0, 2.0]
$t_{\mathrm{age}}$ Age of the cluster in Gyr. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$]
$\tau$ SFH e-folding time in Gyr. Top Hat: [0.1, 10]
@@ 2.3 Source Redshift
In strong lensing, the source redshift is typically constrained either with spectroscopic follow-up (e.g., Sharon et al. 2020) or by inferring a photometric redshift (e.g., Cerny et al. 2018). However, obtaining these redshifts for all identified source galaxies requires prolonged additional inquiry in tension with the appeal of a robust and efficient mass estimator. Remolina González et al. (2020) have shown that substituting a single known source-galaxy redshift with a distribution of source-galaxy redshifts introduces a statistically insignificant uncertainty into the final distribution of mass measurements when compared to the magnitude of other systematic uncertainties. This distribution of lensed source redshifts is measurable (Bayliss et al., 2011a, b; Tran et al., 2022), and we simply adopt a well-described Gaussian with $\mu=2$ and $\sigma=0.2$ from Bayliss et al. (2011a) in keeping with the methodology of Remolina González et al. (2020) as the distribution of source redshifts across the entire sample.
@@ 2.4 Photometry
[…]
@@ 3.1 Total Mass
Following Remolina González et al. (2020), the enclosed total mass within the Einstein aperture was calculated using Equations (1) and (2), where $D(z)$ refers to the angular diameter distance at redshift $z$, via a Monte Carlo approach. The enclosed total mass was computed 1000 times in which each calculation used a different Einstein radius, source redshift, and lens redshift randomly drawn from a Gaussian distribution with standard deviation equal to the uncertainty on each measurement and mean equal to the value for each measurement. Also described in Remolina González et al. (2020) is the need to apply an empirical correction described in Equation (3) based on the “completeness” of the lensed galaxy arcs in each system, where $f(\theta_{E})$ is a cubic polynomial function specified in Table 1 of Remolina González et al. (2020). This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system in which higher Einstein-radius systems have a stronger offset. The correction factor aims to remove this dependence as a function of Einstein radius in non-perfectly-symmetric systems.
@@
For the purposes of applying this correction in each of the 177 lensing systems in this work, we measured the azimuthal coverage ($\phi$) by determining what percentage of the Einstein aperture was traced out by the lensed galaxy arcs in a given system. For example, a lens with a tangential arc stretching from an azimuthal angle of 10 to 100 degrees would have $\phi=0.25$. This was measured by determining what azimuthal fraction of the perimeter of the Einstein aperture was subtended by the the regions masking the source arcs described in Section 2.4. As was shown by Remolina González et al. (2020), whether or not the correction factor was needed was determined by using $\phi$ as an observable proxy for the degree of symmetry for each system. Larger-Einstein-radius systems with a predominantly low $\phi$ tend to deviate from a symmetric dark-matter halo and thus require the empirical correction. For spherically-symmetric systems with a large $\phi$, the measured $M_{\Sigma}(<\theta_{E})$ is taken to be fairly unbiased, and thus the correction factor is not needed. Adopting the convention used by Remolina González et al. 2020 for each system, if $\phi<0.5$, Equation (3) was applied. If $\phi\geq 0.5$, it was not applied. After accounting for the empirical correction, the mean and standard deviation of the resulting 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and error respectively for each system.

### 24 Jan 2024 v3

@@ Abstract
We compute parametric measurements of the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically-selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $0.2\lessapprox z\lessapprox 1.0$ using only two measured parameters in each lensing system: the Einstein radius, and the brightest-cluster-galaxy (BCG) redshift. We then constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions (SEDs) with aperture photometry from the Dark Energy Camera Legacy Survey (DECaLS) in the $g$, $r$, and $z$-band Dark Energy Camera (DECam) filters. We find that the BCG redshift, enclosed total mass, and enclosed luminosity are strongly correlated and well described by a planar relationship in 3D space. We also find that the enclosed total mass and stellar mass are correlated with a logarithmic slope of $0.443\pm 0.035$, and the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope of $-0.563\pm 0.035$. The correlations described here can be used to validate strong-lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time (LSST)—in which an algorithmic treatment of lensing systems will be needed due to the sheer volume of data these surveys will produce.
@@ 1 Introduction
Strong gravitational lensing is a rare phenomenon in the universe in which an intervening object with sufficient surface mass density bends light from a background source relative to the observer such that multiple images of the source are formed. For cases where the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arclike shape (e.g., Lynds & Petrosian 1986; Soucail et al. 1987; Kneib & Natarajan 2011; Meneghetti et al. 2013; Rivera-Thorsen et al. 2017; Shajib et al. 2022). The discovery of new gravitational lenses of this type is accelerating rapidly, particularly with the help of machine learning applied to finding galaxy-galaxy strong lenses in large optical-imaging surveys (e.g., Huang et al. 2021; Rojas et al. 2022; Zaborowski et al. 2023). Coupled with access to a flood of new data from large surveys such as Rubin/LSST in the coming years, the population of candidate lenses both in the bulk and at the margins will only grow. Contemporary methodologies often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (e.g., Diehl et al. 2017); such inspection has been successful in locating new systems thus far (e.g., Rojas et al. 2023). However, with thousands to millions of lines of sight to choose from in an ever-growing plethora of imaging data, statistical measures and correlations in existing lensing systems may help hone candidate-lens samples by quantifying instances of strong lensing in tandem with traditional morphological identification. This would produce a purer sample and help streamline the ability to identify targets. In this paper, we analyze a sample of 177 strong gravitational lenses identified via visual inspection by the COOL-LAMPS collaboration in DECaLS Legacy Survey Data Release 8 (LS DR8, Dey et al. 2019) images, and we investigate the scaling relations between the BCG redshift, Einstein-radius-enclosed (enclosed) total mass, enclosed luminosity, enclosed stellar mass, and enclosed stellar-to-total mass fraction therein. The sample we use here consists of a refined subsample of all strong-lensing candidates identified by COOL-LAMPS, and we consider them unambiguous and robust as informed by their visual morphology—despite a lack of definite spectroscopic confirmation.
@@ 2 Methodology
Galaxies and galaxy clusters host their mass in a variety of forms. Theory and observations both place strong constraints on observed stellar luminosity as a function of mass (e.g., Kuiper 1938; Presotto et al. 2014; Wang & Zhong 2018), hot intracluster gas contributes significantly to the total mass in galaxy clusters (e.g., Forman et al. 1972; Kellogg et al. 1972; Markevitch & Vikhlinin 2007), and both galaxies and galaxy clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al. 1996; Bryan & Norman 1998; Hradecky et al. 2000). However, robustly measuring these mass components on a per-system basis is time and resource intensive—which is unsustainable for large samples. Conveniently, the existence of lensing prescribes a specific amount of mass that must exist to create the observed lensing a priori.
@@
Remolina González et al. (2020) have shown that knowing only the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$), the core mass interior to the Einstein radius of a strong-lensing galaxy cluster may be measured with minimal scatter on order of $\approx 10\%$. With these three parameters, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius. Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties of the lens as well. By directly integrating over pixels inside this aperture in flux-calibrated images, the lens-galaxy photometry can be automatically utilized by an SED modeller to obtain measurements of enclosed luminosity and stellar mass.
@@
The COOL-LAMPS collaboration has constructed a large dataset of strong-lensing candidates using a simple human-led ranking system (e.g., Khullar et al. 2021; Sukay et al. 2022; Martinez et al. 2023; Zhang et al. 2023; Napier et al. 2023b; Klein et al. 2024). From this, we have collated a sample of 177 galaxy clusters and groups that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed source arc is roughly circular with its projected center located at or very near to the BCG. Any systems with complex source-galaxy morphology are less likely to be well described by the simple symmetric dark-matter profiles that inform the parametric mass estimates described in Remolina González et al. (2020)—such complex systems were intentionally excluded from this work.
@@
Figure 1: A subset of nine strong gravitational lenses analyzed in this work. The lenses are arranged by increasing redshift from the top to the bottom of the figure, and they are arranged by increasing Einstein radius from the left to the right of the figure. In each subplot, red circles denote positional constraints used to derive the Einstein radius, white circles represent the Einstein aperture, and a black ‘x’ marks the exact coordinates of the BCG. Respective lens redshifts and Einstein radii are also shown in each subplot. RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using $z$, $r$, and $g$-band imaging data with custom color scaling respectively.
@@
Figure 2: Observed-frame spectral energy distributions (SEDs) in maggies for the nine lenses shown in Figure 1 after SED fitting with Prospector. The top plot of each subplot contains the best-fit SED (red line), 1024 SEDs drawn from their respective posterior distributions (grey lines), observed lens-galaxy photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam, Flaugher et al. 2015) $g$, $r$, and $z$-band filters from left to right respectively (in arbitrary units). The bottom plot of each subplot shows the residual between the observed photometry and the best-fit SED—normalized by the standard deviation of the observed photometry.
@@ 2.1 Einstein Radius
We defined the Einstein radius for each system as the radius of a circle centered on the BCG (Remolina González et al., 2020) which minimizes the total angular separation between each of the three most readily identifiable bright “clumps” in the tangentially lensed arcs and the perimeter of said circle. Henceforth, we refer to this circular aperture centered on the BCG with radius equal to the Einstein radius as the “Einstein aperture”. Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially lensed arc respectively. If three distinct peaks were unable to be found along the tangential arcs, the main arc was simply traced. We bootstrapped the three tangential-arc positional constraints 1000 times, and we adopted the mean and standard deviation from the resulting distribution of Einstein radii as the Einstein radius and its error respectively for each system. A descriptive visualization sampling the range of Einstein radii and lens redshifts considered in this work is shown in Figure 1.
@@ 2.2 Lens Redshift
BCGs are sui generis red-sequence cluster galaxies which can also be used as proxies to infer properties of the cluster-scale dark-matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al. 2014). Since we assume that the lensing systems in this work can be well described by a symmetric dark-matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale dark-matter halo. The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we queried the Sloan Digital Sky Survey Data Release 15 (SDSS DR15, Aguado et al. 2019; Kollmeier et al. 2019) in tandem with the Legacy Survey Data Release 9 (LS DR9, Dey et al. 2019; Duncan 2022) for each system in this work to obtain lens (BCG) redshifts. If the centroid coordinates for each BCG corresponded to a warning-free spectroscopic redshift in SDSS DR15, we adopted that redshift as the lens redshift for that system along with its associated error. If a given BCG lacked any corresponding value in SDSS DR15, as is the case for the higher redshift clusters in this work and/or clusters located in SDSS-non-imaged areas, we adopted the LS DR9 photometric redshift (Zhou et al., 2023) as the lens redshift and its associated error instead. We obtained spectroscopic redshifts for 41 systems and photometric redshifts for the remaining 136 systems.
Table 1: Parameters used in SED Fitting with Prospector
Free Parameter Description Priors
Y $\mathrm{log}(M_{tot}/M_{\odot})$ Total stellar mass formed in dex solar masses. Top Hat: [8.0, 14.0].
Y $\mathrm{log}(Z/Z_{\odot})$ Stellar metallicity in dex solar metallicity. Top Hat: [-1.0, 0.2].
Y $\lambda_{2}$ Diffuse dust optical depth. Top Hat: [0.0, 2.0].
Y $t_{\mathrm{age}}$ Age of the cluster in Gyr. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$].
Y $\tau$ SFH e-folding time in Gyr. Top Hat: [0.1, 10].
N imf_type Initial mass function type. Chabrier (Chabrier, 2003).
N dust_type Dust attenuation curve. Calzetti (Calzetti et al., 2000).
N sfh Star formation history model. Delayed tau (Carnall et al., 2019).
@@ 2.3 Source Redshift
In strong lensing, the source redshift is typically constrained either with spectroscopic follow-up (e.g., Sharon et al. 2020) or by inferring a photometric redshift (e.g., Cerny et al. 2018). However, obtaining these redshifts for all identified source galaxies in a given system requires prolonged additional inquiry in tension with the appeal of an efficient mass estimator. Remolina González et al. (2020) have shown that substituting a single known source-galaxy redshift with a distribution of source-galaxy redshifts introduces a statistically insignificant uncertainty into the final distribution of mass measurements when compared to the magnitude of other systematic uncertainties. This distribution of lensed source redshifts is measurable (e.g., Bayliss et al. 2011a, b; Tran et al. 2022), and we simply adopt a well-described Gaussian with $\mu=2$ and $\sigma=0.2$ from Bayliss et al. (2011a) in keeping with the methodology of Remolina González et al. (2020) as the distribution of source redshifts across the entire sample in this work.
@@ 2.4 Photometry
We obtained aperture photometry for SED fitting in each system by linearly summing the dereddened flux from pixels within the Einstein aperture in $g$, $r$, and $z$-band imaging data from DECaLS LS DR9. Since it is critical that we do not include light from the lensed source arcs as well as any intervening stars or foreground galaxies in the measured photometry, we created two masks to exclude non-lens-galaxy regions. The first mask removed flux from the lensed source arcs, and the second mask removed flux from interlopers as identified by a color and magnitude screening by eye. Accounting for masking, this linear integration was done for each of the 1000 bootstrapped Einstein apertures derived in Section 2.1, and the resulting mean and standard deviation were adopted as the $g$, $r$, and $z$-band photometry and their errors respectively for each system.
@@ 3.1 Total Mass
Following Remolina González et al. (2020), the enclosed total mass within the Einstein aperture was calculated using Equations (1) and (2)—where $D(z)$ refers to the angular diameter distance at redshift $z$—via a Monte Carlo approach. The enclosed total mass was computed 1000 times for each system, in which each calculation used a different Einstein radius, source redshift, and lens redshift randomly drawn from a Gaussian distribution with mean equal to the parameter value and standard deviation equal to the parameter error. Also described in Remolina González et al. (2020) is the need to apply an empirical correction described in Equation (3) based on the “completeness” of the lensed source arcs in each system—where $f(\theta_{E})$ is a cubic polynomial function specified in Table 1 of Remolina González et al. (2020). This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system, in which larger-Einstein-radius systems have a stronger offset. The correction factor aims to remove this dependence as a function of Einstein radius in non-perfectly symmetric systems.
@@
For the purposes of applying this correction in each of the 177 lensing systems analyzed in this work, we measured the azimuthal coverage ($\phi$)—defined as the percentage of the Einstein aperture that was traced out by the lensed source arcs in a given system. For example, a lens with a tangential arc stretching from an azimuthal angle of 10 to 100 degrees would have $\phi=0.25$. As was done by Remolina González et al. (2020), $\phi$ was used as an observable proxy for the degree of symmetry in each system to determine whether or not the correction factor was needed. Larger-Einstein-radius systems with a predominantly low $\phi$ tend to deviate from a symmetric dark-matter halo and thus require the empirical correction. For spherically symmetric systems with a large $\phi$, the measured $M_{\Sigma}(<\theta_{E})$ is taken to be fairly unbiased, and thus the correction factor is not needed. We obtained $\phi$ by determining the fraction of the perimeter of the Einstein aperture that was subtended by the the regions masking the lensed source arcs described in Section 2.4 in each system. Adopting the convention used by Remolina González et al. (2020) in their analysis, if $\phi<0.5$, Equation (3) was applied. If $\phi\geq 0.5$, it was not applied. Accounting for the empirical correction, the mean and standard deviation of the 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and its error respectively for each system.